In the Dicovering Algebra: An Investigative approach, there is a question that says Find the value of the unknown number in each proportion? The problem is 52/91=42/s. So do i inverse it. Swich it up like this: 91/52=s/42? or keep it the same?
Answer by Alan3354(69443)

Find the value of the unknown number in each proportion? The problem is 52/91=42/s. So do i inverse it. Swich it up like this: 91/52=s/42? or keep it the same?
----------------------
91/52=42/s
You want to get s by itself on one side of the equal sign.
You don't invert anything in equations.
A shortcut is to cross multiply. Multiply each NUM by the DEN of the other fraction.
91s = 52*42
Then divide by 91
s = 52*42/91
Then reduce the fraction
s = 4*42/7
s = 4*6
s = 24